CLEO PUBLICATIONS

March 20, 2008

1. Observation of Three Upsilon States D. Andrews et al., Physical Review Letters 44, 1108 (1980)

2. Observation of a Fourth Upsilon State in e^{+}e^{−} Annihilations D. Andrews et al., Physical Review Letters 45, 219 (1980)

3. Evidence for New-Flavor Production at the Υ(4S)

C. Bebek et al., Physical Review Letters 46, 84 (1981)

4. Decay of b-flavored Hadrons to Single-Muon and Dimuon Final States K. Chadwick et al., Physical Review Letters 46, 88 (1981)

5. Measurement of the Branching of Υ(2S) → π^{+}π^{−} + Υ(1S)

J.J. Mueller et al., Physical Review Letters 46, 1181 (1981)

6. Decay of B Mesons into Charged and Neutral Kaons A. Brody et al., Physical Review Letters 48, 1070 (1982)

7. Charged-Particle Multiplicities in B-Meson Decay M.S. Alam et al., Physical Review Letters 49, 357 (1982)

8. Inclusive Charged D^{∗} Production in e^{+}e^{−} Annihilations at W = 10.4 GeV C. Bebek et al., Physical Review Letters 49, 610 (1982)

9. Measurement of the Branching Ratio of Υ(3S) → π^{+}π^{−}Υ(1S)

J. Green et al., Physical Review Letters 49, 617 (1982)

10. Semileptonic Decay of B Mesons K. Chadwick et al., Physical Review D 27, 475 (1983)

11. Ruling out Exotic Models of b Quark Decay A. Chen et al., Physics Letters B 122, 317 (1983)

12. Measurement of the Leptonic Branching Ratios of the Υ(1S), Υ(2S), and Υ(3S) D. Andrews et al., Physical Review Letters 50, 807 (1983)

13. Measurement of the Branching Fraction of the Decay Υ(1S) → τ^{+}τ^{−} R. Giles et al., Physical Review Letters 50, 877 (1983)

14. Observation of Exclusive Decay Modes of b-flavored Mesons S. Behrends et al., Physical Review Letters 50, 881 (1983)

15. Search for Axion Production in Υ Decay M.S. Alam et al., Physical Review D 27, 1665 (1983)

16. The CLEO Detector D. Andrews et al., Nuclear Instruments and Methods 211, 47 (1983)

17. D^{0} Spectrum from B-Meson Decay J. Green et al., Physical Review Letters 51, 347 (1983)

18. Evidence for the F Meson at 1970 MeV A. Chen et al., Physical Review Letters 51, 634 (1983)

19. Production of Charmed Mesons in e^{+}e^{−} Annihilation at 10.5 GeV P. Avery et al., Physical Review Letters 51, 1139 (1983)

20. Observation of Baryons in B-Meson Decay M.S. Alam et al., Physical Review Letters 51, 1143 (1983)

21. Observation of Radiative Decays of the Υ(2S)

P. Haas et al., Physical Review Letters 52, 799 (1984)

22. Limit on the b → u Coupling from Semileptonic B Decay A. Chen et al., Physical Review Letters 52, 1084 (1984)

23. A Search for the ξ(2.2) in the Υ Region S. Behrends et al., Physics Letters B 137, 277 (1984)

24. Total Cross Section for Electron-Positron Annihilation into Hadron Final States in the Upsilon Energy Region R. Giles et al., Physical Review D 29, 1285 (1984)

25. Hyperon Production in e^{+}e^{−} Interactions in the Υ Region M.S. Alam et al., Physical Review Letters 53, 24 (1984)

26. Upper Limit on Flavor-Changing Neutral-Current Decays of the b Quark P. Avery et al., Physical Review Letters 53, 1309 (1984)

27. High-Statistics Study of Υ(2S) → π^{+}π^{−}Υ(1S)

D. Besson et al., Physical Review D 30, 1433 (1984)

28. Leptonic Branching Ratio of the Υ(2S)

P. Haas et al., Physical Review D 30, 1996 (1984)

29. Two-Body Decays of B Mesons R. Giles et al., Physical Review D 30, 2279 (1984)

30. Observation of New Structure in the e^{+}e^{−} Cross Section above the Υ(4S) D. Besson et al., Physical Review Letters 54, 381 (1985)

31. Inclusive Decay of B Mesons into Charged D^{∗} S.E. Csorna et al., Physical Review Letters 54, 1894 (1985)

32. Inclusive Hadron Production in Upsilon Decays and in Nonresonant Electron–Positron Annihilation at 10.49 GeV S. Behrends et al., Physical Review D 31, 2161 (1985)

33. Observation of the decay \overline{B}^{0} → D^{∗+}ρ^{−} A. Chen et al., Physical Review D 31, 2386 (1985)

34. Λ_{c} Production from e^{+}e^{−} Annihilation in the Υ Energy Region T. Bowcock et al., Physical Review Letters 55, 923 (1985)

35. Decay B → ψX P. Haas et al., Physical Review Letters 55, 1248 (1985)

36. Bose-Einstein Correlations in e^{+}e^{−} Annihilations in the Υ Region P. Avery et al., Physical Review D 32, 2294 (1985)

37. New Determination of the Michel Parameter in Tau Decay S. Behrends et al., Physical Review D 32, 2468 (1985)

38. Search for Monoenergetic Photons from Υ(1S) → γ + X D. Besson et al., Physical Review D 33, 300 (1986)

39. Inclusive ϕ production in B-Meson Decay D. Bortoletto et al., Physical Review Letters 56, 800 (1986)

40. Measurement of the Direct Photon Spectrum from the Υ(1S)

S.E. Csorna et al., Physical Review Letters 56, 1222 (1986)

41. Decay D^{0} → ϕ\overline{K}^{0} C. Bebek et al., Physical Review Letters 56, 1893 (1986)

42. Upper Limits for the Production of Light Short-Lived Neutral Particles in Radiative Υ Decay T. Bowcock et al., Physical Review Letters 56, 2676 (1986)

43. Observation of the Decay B → F X P. Haas et al., Physical Review Letters 56, 2781 (1986)

44. Exclusive Radiative Υ(1S) Decays A. Bean et al., Physical Review D 34, 905 (1986)

45. Study of the Decay B → ψX M.S. Alam et al., Physical Review D 34, 3279 (1986)

46. Inclusive B-Meson Decays into Charm D. Bortoletto et al., Physical Review D 35, 19 (1987)

47. Limits on Rare Exclusive Decays of B Mesons P. Avery et al., Physics Letters B 183, 429 (1987)

48. Limits on B^{0}\overline{B}^{0} Mixing and τ_{B^{0}}/τ_{B^{+}} A. Bean et al., Physical Review Letters 58, 183 (1987)

49. Study of π^{+}π^{−} transitions from the Υ(3S)

T. Bowcock et al., Physical Review Letters 58, 307 (1987)

50. Search for Magnetically Charged Particles Produced in e^{+}e^{−} Annihilations at \sqrt{s} = 10.6 GeV T. Gentile et al., Physical Review D 35, 1081 (1987)

51. Limit on the Mass of the Tau Neutrino S.E. Csorna et al., Physical Review D 35, 2747 (1987)

52. Branching Ratios of B Mesons to K^{+}, K^{−}, and K^{0}/\overline{K}^{0} M.S. Alam et al., Physical Review Letters 58, 1814 (1987)

53. Improved Upper limit on Flavor-Changing Neutral-Current Decays of the b Quark A. Bean et al., Physical Review D 35, 3533 (1987)

54. Measurement of the D^{0}, D^{+}, and D_{s}^{+} Meson Lifetimes at 10.58 GeV S.E. Csorna et al., Physics Letters B 191, 318 (1987)

55. Evidence for Charmed Baryons in B-Meson Decay M.S. Alam et al., Physical Review Letters 59, 22 (1987)

56. Γ(b → uℓν)/Γ(b → cℓν) from the End Point of the Lepton Momentum Spectrum in Semileptonic B Decay S. Behrends et al., Physical Review Letters 59, 407 (1987)

57. Measurement of the Tau Lifetime C. Bebek et al., Physical Review D 36, 690 (1987)

58. Exclusive Decays and Masses of the B Mesons C. Bebek et al., Physical Review D 36, 1289 (1987)

59. Production of η and ω Mesons in τ Decay and a Search for Second-Class Currents P. Baringer et al., Physical Review Letters 59, 1993 (1987)

60. Charm Production in Nonresonant e^{+}e^{−} Annihilations at \sqrt{s}=10.55 GeV D. Bortoletto et al., Physical Review D 37, 1719 (1988) Erratum Physical Review D 39, 1471 (1989)

61. Upper Limits on Charm-Changing Neutral-Current Interactions P. Haas et al., Physical Review Letters 60, 1614 (1988)

62. Investigation of the Total Charm-Pair Cross Section in Nonresonant e^{+}e^{−} Annihilations at \sqrt{s} = 10.5 GeV T. Bowcock et al., Physical Review D 38, 2679 (1988) Erratum Physical Review D 40, 1701 (1989)

63. Search for Charmless Decays B → p\overline{p}π and B → p\overline{p}ππ C. Bebek et al., Physical Review Letters 62, 8 (1989)

64. Observation of the Charmed Strange Baryon Ξ_{c}^{0} P. Avery et al., Physical Review Letters 62, 863 (1989)

65. Σ_{c}^{++} and Σ_{c}^{0} Production from e^{+}e^{−} Annihilation in the Υ Energy Region T. Bowcock et al., Physical Review Letters 62, 1240 (1989)

66. B^{0}\overline{B}^{0} mixing at the Υ(4S)

M. Artuso et al., Physical Review Letters 62, 2233 (1989)

67. Search for b → u Transitions in Exclusive Hadronic B-Meson Decays D. Bortoletto et al., Physical Review Letters 62, 2436 (1989)

68. Measurement of the Muonic Branching Fractions of the Υ(1S) and the Υ(3S) W.-Y. Chen et al., Physical Review D 39, 3528 (1989)

69. A Search for Exclusive Penguin Decays of B Mesons P. Avery et al., Physics Letters B 223, 470 (1989)

70. Search for the Production of Fractionally Charged Particles in e^{+}e^{−} Annihilations at \sqrt{s} = 10.5 GeV T. Bowcock et al., Physical Review D 40, 263 (1989)

71. First Observation of Inclusive ψ Production in Υ Decays R. Fulton et al., Physics Letters B 224, 445 (1989)

72. Search for a Neutral Higgs Boson in B-Meson Decay M.S. Alam et al., Physical Review D 40, 712 (1989) Erratum Physical Review D 40, 3790 (1989)

73. Measurement of D_{s} Decay Modes W.-Y. Chen et al., Physics Letters B 226, 192 (1989)

74. Measurement of the Isospin Mass Splitting Ξ_{c}^{+} – Ξ_{c}^{0} M.S. Alam et al., Physics Letters B 226, 401 (1989)

75. Study of the Decay \overline{B}^{0} → D^{∗+}l^{−}\overline{ν} D. Bortoletto et al., Physical Review Letters 63, 1667 (1989)

76. Observation of B-Meson Semileptonic Decays to Noncharmed Final States R. Fulton et al., Physical Review Letters 64, 16 (1990)

77. P-Wave Charmed Mesons in e^{+}e^{−} Annihilation P. Avery et al., Physical Review D 41, 774 (1990)

78. Search for Neutrinoless Decays of the τ Lepton T. Bowcock et al., Physical Review D 41, 805 (1990)

79. Radiative Υ(1S) Decays R. Fulton et al., Physical Review D 41, 1401 (1990)

80. Exclusive and Inclusive Decays of B Mesons into D_{s} Mesons D. Bortoletto et al., Physical Review Letters 64, 2117 (1990)

81. Observation of Υ(4S) Decays into non-B\overline{B} Final States Containing ψ Mesons J.P. Alexander et al., Physical Review Letters 64, 2226 (1990)

82. Measurement of γγ Widths of Charmonium States W.-Y. Chen et al., Physics Letters B 243, 169 (1990)

83. Study of the Decays D^{0} → K\overline{K}, π\overline{π} J.P. Alexander et al., Physical Review Letters 65, 1184 (1990)

84. Determination of B(D_{s}^{+} → ϕπ^{+}) via Observation of D_{s}^{+} → ϕℓ^{+}ν J.P. Alexander et al., Physical Review Letters 65, 1531 (1990)

85. Study of K^{∗} Production in Tau Decay M. Goldberg et al., Physics Letters B 251, 223 (1990)

86. Measurement of the Λ_{c}^{+} Decay-Asymmetry Parameter P. Avery et al., Physical Review Letters 65, 2842 (1990)

87. Exclusive and Inclusive Semileptonic Decays of B Mesons to D Mesons R. Fulton et al., Physical Review D 43, 651 (1991)

88. Study of π^{+}π^{−} Transitions from the Υ(3S) and a Search for the h_{b} I.C. Brock et al., Physical Review D 43, 1448 (1991)

89. Study of D^{0} Decays into Final States with a π^{0} or η K. Kinoshita et al., Physical Review D 43, 2836 (1991)

90. Inclusive Production of the Charmed Baryon Λ_{c}^{+} from e^{+}e^{−} Annihilations at \sqrt{s} = 10.55 GeV P. Avery et al., Physical Review D 43, 3599 (1991)

91. Study of Continuum D^{∗+} Spin Alignment Y. Kubota et al., Physical Review D 44, 593 (1991)

92. Measurement of the Inclusive B^{∗} Cross Section above the Υ(4S)

D.S. Akerib et al., Physical Review Letters 67, 1692 (1991)

93. Inclusive χ(2P) Production in Υ(3S) Decay R. Morrison et al., Physical Review Letters 67, 1696 (1991)

94. Unusual Decay Modes of D^{0} and D^{+} Mesons R. Ammar et al., Physical Review D 44, 3383 (1991)

95. Measurement of the Ratio B(D^{0} → K^{∗−}e^{+}ν_{e})/B(D^{0} → K^{−}e^{+}ν_{e})

G. Crawford et al., Physical Review D 44, 3394 (1991)

96. Inclusive and Exclusive Decays of B Mesons to Final States Including Charm and Charmonium Mesons D. Bortoletto et al., Physical Review D 45, 21 (1992)

97. Measurement of Baryon Production in B-Meson Decay G. Crawford et al., Physical Review D 45, 752 (1992)

98. Measurements of Semileptonic Branching Fractions of B Mesons at the Υ(4S) Resonance S. Henderson et al., Physical Review D 45, 2212 (1992)

99. Two-Body D_{s}^{+} decays to ηπ^{+}, η′π^{+}, ηρ^{+}, η′ρ^{+}, and ϕρ^{+} M. Daoudi et al., Physical Review D 45, 3965 (1992)

100. D_{s}^{+} Decays to ηπ^{+} and η′π^{+} J.P. Alexander et al., Physical Review Letters 68, 1275 (1992)

101. D_{s}^{+} decays to ηρ^{+}, η′ρ^{+}, and ϕρ^{+} P. Avery et al., Physical Review Letters 68, 1279 (1992)

102. Electronic Branching Ratio of the τ Lepton R. Ammar et al., Physical Review D 45, 3976 (1992)

103. Observation of the Decay Ξ_{c}^{0} → Ω^{−}K^{+} S. Henderson et al., Physics Letters B 283, 161 (1992)

104. The CLEO II Detector Y. Kubota et al., Nuclear Instruments and Methods A 320, 66 (1992)

105. A Measurement of the Tau Lepton Lifetime M. Battle et al., Physics Letters B 291, 488 (1992)

106. Measurement of the D^{∗}(2010) Branching Fractions F. Butler et al., Physical Review Letters 69, 2041 (1992)

107. Isospin Mass Splittings from Precision Measurements of D^{∗} − D Mass Differences D. Bortoletto et al., Physical Review Letters 69, 2046 (1992)

108. Exclusive χ(2P) Production in Υ(3S) Decay G. Crawford et al., Physics Letters B 294, 139 (1992)

109. Shape Studies of Quark Jets vs. Gluon Jets at \sqrt{s} = 10 GeV M.S. Alam et al., Physical Review D 46, 4822 (1992)

110. Measurement of τ Decays Involving η Mesons M. Artuso et al., Physical Review Letters 69, 3278 (1992)

111. Measurement of the Tau Lepton Electronic Branching Fraction D.S. Akerib et al., Physical Review Letters 69, 3610 (1992) Erratum Physical Review Letters 71, 3395 (1993)

112. Search for τ^{−} → γμ^{−}: A Test of Lepton Number Conservation A. Bean et al., Physical Review Letters 70, 138 (1993)

113. Lepton Asymmetry Measurements in \overline{B} → D^{∗}l^{−}\overline{ν}_{l} and Implications for V − A and the Form Factors S. Sanghera et al., Physical Review D 47, 791 (1993)

114. Tau Decays with One Charged Particle plus Multiple π^{0}’s M. Procario et al., Physical Review Letters 70, 1207 (1993)

115. Production and Decay of the D_{s1}(2536)^{+} J.P. Alexander et al., Physics Letters B 303, 377 (1993)

116. Search for Exclusive b → u Semileptonic Decays of B Mesons A. Bean et al., Physical Review Letters 70, 2681 (1993)

117. Measurement of the τ-Lepton Mass R. Balest et al., Physical Review D 47, R3671 (1993)

118. Limit on the Tau Neutrino Mass D. Cinabro et al., Physical Review Letters 70, 3700 (1993)

119. Evidence for Penguin-Diagram Decays: First Observation of B → K^{∗}(892)γ R. Ammar et al., Physical Review Letters 71, 674 (1993)

120. Measurement of the Ratio B(D^{+} → π^{0}ℓ^{+}ν)/B(D^{+} → \overline{K}^{0}ℓ^{+}ν)

M.S. Alam et al., Physical Review Letters 71, 1311 (1993)

121. Two Measurements of B^{0}\overline{B}^{0} Mixing J. Bartelt et al., Physical Review Letters 71, 1680 (1993)

122. Measurement of the Decay τ^{−} → π^{−}π^{+}π^{−}2π^{0}ν_{τ} D. Bortoletto et al., Physical Review Letters 71, 1791 (1993)

123. Measurement of the D → ππ Branching Fractions M. Selen et al., Physical Review Letters 71, 1973 (1993)

124. Study of the Decays Λ_{c}^{+} → Ξ^{0}K^{+}, Λ_{c}^{+} → Σ^{+}K^{+}K^{−} and Λ_{c}^{+} → Ξ^{−}K^{+}π^{+} P. Avery et al., Physical Review Letters 71, 2391 (1993)

125. Study of D^{0} decays into \overline{K}^{0} and \overline{K}^{∗0} M. Procario et al., Physical Review D 48, 4007 (1993)

126. Measurement of the Absolute Branching Fraction for D^{0} → K^{−}π^{+} D.S. Akerib et al., Physical Review Letters 71, 3070 (1993)

127. Measurement of Exclusive Λ_{c} Decays with a Σ^{+} in the Final State Y. Kubota et al., Physical Review Letters 71, 3255 (1993)

128. Observation of the Charmed Baryon Σ_{c}^{+} and Measurement of the Isospin Mass Splittings of the Σ_{c} G. Crawford et al., Physical Review Letters 71, 3259 (1993)

129. Measurements of Exclusive Semileptonic Decays of D Mesons A. Bean et al., Physics Letters B 317, 647 (1993)

130. Observation of B^{0} Decay to Two Charmless Mesons M. Battle et al., Physical Review Letters 71, 3922 (1993)

131. Measurement of Charmless Semileptonic Decays of B Mesons J. Bartelt et al., Physical Review Letters 71, 4111 (1993)

132. Search for Exclusive b → u Transitions in Hadronic Decays of B Mesons Involving D_{s}^{+} and D_{s}^{∗+} Mesons J.P. Alexander et al., Physics Letters B 319, 365 (1993)

133. Analysis of Hadronic Transitions in Υ(3S) Decays F. Butler et al., Physical Review D 49, 40 (1994)

134. Study of the Decay Λ_{c}^{+} → Λℓ^{+}ν_{ℓ} T. Bergfeld et al., Physics Letters B 323, 219 (1994)

135. Observation of D^{0} → K^{+}π^{−} D. Cinabro et al., Physical Review Letters 72, 1406 (1994)

136. Observation of a New Charmed Strange Meson Y. Kubota et al., Physical Review Letters 72, 1972 (1994)

137. Measurement of the Branching Fraction for D^{+} → K^{−}π^{+}π^{+} R. Balest et al., Physical Review Letters 72, 2328 (1994)

138. A Measurement of B(D_{s}^{+} → ϕℓ^{+}ν_{ℓ})/B(D_{s}^{+} → ϕπ^{+})

F. Butler et al., Physics Letters B 324, 255 (1994)

139. Observation of Λ_{c}^{+} Decays to Λπ^{+}π^{0}, Σ^{0}π^{+}, Σ^{0}π^{+}π^{0}, and Σ^{0}π^{−}π^{+}π^{+} P. Avery et al., Physics Letters B 325, 257 (1994)

140. First Measurement of Γ(D_{s}^{+} → μ^{+}ν)/Γ(D_{s}^{+} → ϕπ^{+})

D. Acosta et al., Physical Review D 49, 5690 (1994)

141. Search for B^{0} Decays to Two Charged Leptons R. Ammar et al., Physical Review D 49, 5701 (1994)

142. A Measurement of the Branching Fraction B(τ^{−} → h^{−}π^{0}ν_{τ})

M. Artuso et al., Physical Review Letters 72, 3762 (1994)

143. Production and Decay of D_{1}(2420)^{0} and D_{2}^{∗}(2460)^{0} P. Avery et al., Physics Letters B 331, 236 (1994) Erratum Physics Letters B 342, 453 (1995)

144. Exclusive Hadronic B Decays to Charm and Charmonium Final States M.S. Alam et al., Physical Review D 50, 43 (1994)

145. Precision Measurement of the D_{s}^{∗+} – D_{s}^{+} Mass Difference D.N. Brown et al., Physical Review D 50, 1884 (1994)

146. Two-Photon Production of Charged Pion and Kaon Pairs J. Dominick et al., Physical Review D 50, 3027 (1994)

147. Measurement of Two-Photon Production of the χ_{c2} J. Dominick et al., Physical Review D 50, 4265 (1994)

148. Measurement of the Cross-Section for γγ → p\overline{p} M. Artuso et al., Physical Review D 50, 5484 (1994)

149. Luminosity Measurement with the CLEO II Detector G. Crawford et al., Nuclear Instruments and Methods A 345, 429 (1994)

150. Study of the Five-Charged-Pion Decay of the τ Lepton D. Gibaut et al., Physical Review Letters 73, 934 (1994)

151. Measurement of Cabibbo-suppressed Decays of the τ Lepton M. Battle et al., Physical Review Letters 73, 1079 (1994)

152. Observation of Inclusive B Decays to the Charmed Baryons Σ_{c}^{++} and Σ_{c}^{0} M. Procario et al., Physical Review Letters 73, 1472 (1994)

153. Search for Neutrinoless Decays of the Tau Lepton J. Bartelt et al., Physical Review Letters 73, 1890 (1994)

154. Semileptonic Branching Fractions of Charged and Neutral B Mesons M. Athanas et al., Physical Review Letters 73, 3503 (1994) Erratum Physical Review Letters 74, 3090 (1995)

155. Measurement of the Ratios of Form Factors in the Decay D_{s}^{+} → ϕe^{+}ν_{e} P. Avery et al., Physics Letters B 337, 405 (1994)

156. Measurement of the Branching Fraction for Υ(1S) → τ^{+}τ^{−} D. Cinabro et al., Physics Letters B 340, 129 (1994)

157. Observation of D_{1}(2420)^{+} and D_{2}^{∗}(2460)^{+} T. Bergfeld et al., Physics Letters B 340, 194 (1994)

158. Observation of B → ψπ Decays J.P. Alexander et al., Physics Letters B 341, 435 (1995) Erratum Physics Letters B 347, 469 (1995)

159. Measurement of the \overline{B} → D^{∗}ℓ\overline{ν} Branching Fractions and |V_{cb}| B. Barish et al., Physical Review D 51, 1014 (1995)

160. Υ(1S) → γ + Non-interacting Particles R. Balest et al., Physical Review D 51, 2053 (1995)

161. First Measurement of the Rate for the Inclusive Radiative Penguin Decay b → sγ M.S. Alam et al., Physical Review Letters 74, 2885 (1995)

162. First Observation of the Decay Ξ_{c}^{+} → Ξ^{0}e^{+}ν_{e} and an Estimate of the Ξ_{c}^{+}/Ξ_{c}^{0} Lifetime Ratio J.P. Alexander et al., Physical Review Letters 74, 3113 (1995) Erratum Physical Review Letters 75, 4155 (1995)

163. Observation of Excited Charmed Baryon States Decaying to Λ_{c}^{+}π^{+}π^{−} K.W. Edwards et al., Physical Review Letters 74, 3331 (1995)

164. New Decay Modes of the Λ_{c}^{+} Charmed Baryon R. Ammar et al., Physical Review Letters 74, 3534 (1995)

165. Form Factor Ratio Measurement in Λ_{c}^{+} → Λe^{+}ν_{e} G. Crawford et al., Physical Review Letters 75, 624 (1995)

166. Search for B → ℓ\overline{ν}_{ℓ} M. Artuso et al., Physical Review Letters 75, 785 (1995)

167. Observation of the Isospin-Violating Decay D_{s}^{∗+} → D_{s}^{+}π^{0} J. Gronberg et al., Physical Review Letters 75, 3232 (1995)

168. Measurement of the D_{s}^{+} → ηℓ^{+}ν and D_{s}^{+} → η′ℓ^{+}ν Branching Ratios G. Brandenburg et al., Physical Review Letters 75, 3804 (1995)

169. Measurements of the Decays τ^{−} → h^{−}h^{+}h^{−}ν_{τ} and τ^{−} → h^{−}h^{+}h^{−}π^{0}ν_{τ} R. Balest et al., Physical Review Letters 75, 3809 (1995)

170. Observation of a Narrow State Decaying into Ξ_{c}^{+}π^{−} P. Avery et al., Physical Review Letters 75, 4364 (1995)

171. Measurement of the Decay Asymmetry Parameters in Λ_{c}^{+} → Λπ^{+} and Λ_{c}^{+} → Σ^{+}π^{0} M. Bishai et al., Physics Letters B 350, 256 (1995)

172. Measurement of α_{s} from τ Decays T. Coan et al., Physics Letters B 356, 580 (1995)

173. Measurement of the Ratio of Branching Fractions B(D^{0} → π^{−}e^{+}ν_{e})/B(D^{0} → K^{−}e^{+}ν_{e}) F. Butler et al., Physical Review D 52, 2656 (1995)

174. Inclusive Decays of B Mesons to Charmonium R. Balest et al., Physical Review D 52, 2661 (1995)

175. Search for CP Violation in D^{0} Decay J. Bartelt et al., Physical Review D 52, 4860 (1995)

176. Measurement of the B Semileptonic Branching Fraction with Lepton Tags B. Barish et al., Physical Review Letters 76, 1570 (1996)

177. Tau Decays into Three Charged Leptons and Two Neutrinos M.S. Alam et al., Physical Review Letters 76, 2637 (1996)

178. Limits on Flavor Changing Neutral Currents in D^{0} Meson Decays A. Freyberger et al., Physical Review Letters 76, 3065 (1996) Erratum Physical Review Letters 77, 2147 (1996)

179. Measurement of the Form Factors for \overline{B}^{0} → D^{∗+}ℓ^{−}\overline{ν} J.E. Duboscq et al., Physical Review Letters 76, 3898 (1996)

180. First Observation of the Decay τ^{−} → K^{−}ην_{τ} J. Bartelt et al., Physical Review Letters 76, 4119 (1996)

181. Observation of an Excited Charmed Baryon Decaying into Ξ_{c}^{0}π^{+} L. Gibbons et al., Physical Review Letters 77, 810 (1996)

182. A Search for Nonresonant B^{+} → h^{+}h^{−}h^{+} Decays T. Bergfeld et al., Physical Review Letters 77, 4503 (1996)

183. First Measurement of the B → πℓν and B → ρ(ω)ℓν Branching Fractions J.P. Alexander et al., Physical Review Letters 77, 5000 (1996)

184. Observation of the Ξ_{c}^{+} Charmed Baryon Decays to Σ^{+}K^{−}π^{+}, Σ^{+}\overline{K}^{∗0}, and ΛK^{−}π^{+}π^{+} T. Bergfeld et al., Physics Letters B 365, 431 (1996)

185. Study of B → ψρ M. Bishai et al., Physics Letters B 369, 186 (1996)

186. Observation of New Decay Modes of the Charmed-Strange Baryon Ξ_{c}^{+} K.W. Edwards et al., Physics Letters B 373, 261 (1996)

187. A Measurement of B(D^{0} → K^{−}π^{+}π^{0})/B(D^{0} → K^{−}π^{+})

B. Barish et al., Physics Letters B 373, 334 (1996)

188. Measurement of the Branching Fraction for D_{s}^{−} → ϕπ^{−} M. Artuso et al., Physics Letters B 378, 364 (1996)

189. Measurement of the Tau Lepton Lifetime R. Balest et al., Physics Letters B 388, 402 (1996)

190. Observation of the Cabibbo-Suppressed Charmed Baryon Decay Λ_{c}^{+} → pϕ J.P. Alexander et al., Physical Review D 53, R1013 (1996)

191. Search for Exclusive Charmless Hadronic B Decays D.M. Asner et al., Physical Review D 53, 1039 (1996)

192. Measurements of B → D_{s}^{+}X Decays D. Gibaut et al., Physical Review D 53, 4734 (1996)

193. The Inclusive Decay B → ηX Y. Kubota et al., Physical Review D 53, 6033 (1996)

194. Decays of Tau Leptons to Final States Containing K_{S}^{0} Mesons T.E. Coan et al., Physical Review D 53, 6037 (1996)

195. Measurement of the Inclusive Semielectronic D^{0} Branching Fraction Y. Kubota et al., Physical Review D 54, 2994 (1996)

196. Analysis of D^{0} → K\overline{K}X Decays D.M. Asner et al., Physical Review D 54, 4211 (1996)

197. Observation of Two Excited Charmed Baryons Decaying into Λ_{c}^{+}π^{±} G. Brandenburg et al., Physical Review Letters 78, 2304 (1997)

198. Analyses of D^{+} → K_{S}^{0}K^{+} and D^{+} → K_{S}^{0}π^{+} M. Bishai et al., Physical Review Letters 78, 3261 (1997)

199. A Measurement of the Michel Parameters in Leptonic Decays of the Tau R. Ammar et al., Physical Review Letters 78, 4686 (1997)

200. Search for the Decays B^{0} → D^{(∗)+}D^{(∗)−} D.M. Asner et al., Physical Review Letters 79, 799 (1997)

201. Search for Neutrinoless τ Decays Involving π^{0} or η Mesons G. Bonvicini et al., Physical Review Letters 79, 1221 (1997)

202. Observation of the Decay D_{s}^{+} → ωπ^{+} R. Balest et al., Physical Review Letters 79, 1436 (1997)

203. Measurement of the \overline{B} → Dℓ\overline{ν} Partial Width and Form Factor Parameters M. Athanas et al., Physical Review Letters 79, 2208 (1997)

204. First Observation of τ → 3πην_{τ} and τ → f_{1}πν_{τ} Decays T. Bergfeld et al., Physical Review Letters 79, 2406 (1997)

205. Observation of Exclusive B Decays to Final States Containing a Charmed Baryon X. Fu et al., Physical Review Letters 79, 3125 (1997)

206. First Observation of Inclusive B Decays to the Charmed Strange Baryons Ξ_{c}^{0} and Ξ_{c}^{+} B. Barish et al., Physical Review Letters 79, 3599 (1997)

207. Study of the Decay τ^{−} → 2π^{−}π^{+}3π^{0}ν_{τ} S. Anderson et al., Physical Review Letters 79, 3814 (1997)

208. Limit on the Two-Photon Production of the Glueball Candidate f_{J}(2220) at the Cornell Electron Storage Ring R. Godang et al., Physical Review Letters 79, 3829 (1997)

209. Measurement of the Decay Amplitudes and Branching Fractions of B → J/ψK^{∗} and B → J/ψK Decays C.P. Jessop et al., Physical Review Letters 79, 4533 (1997)

210. Study of the B^{0} Semileptonic Decay Spectrum at the Υ(4S) Resonance M. Artuso et al., Physics Letters B 399, 321 (1997)

211. Studies of the Cabibbo-Suppressed Decays D^{+} → π^{0}ℓ^{+}ν and D^{+} → ηe^{+}ν_{e} J. Bartelt et al., Physics Letters B 405, 373 (1997)

212. Study of Flavor-Tagged Baryon Production in B Decay R. Ammar et al., Physical Review D 55, 13 (1997)

213. Search for ϕ Mesons in τ Lepton Decay P. Avery et al., Physical Review D 55, R1119 (1997)

214. Experimental Tests of Lepton Universality in τ Decay A. Anastassov et al., Physical Review D 55, 2559 (1997)

215. Search for Neutrinoless τ Decays: τ → eγ and τ → μγ K.W. Edwards et al., Physical Review D 55, R3919 (1997)

216. Measurement of the Direct Photon Spectrum in Υ(1S) Decays B. Nemati et al., Physical Review D 55, 5273 (1997)

217. ν_{τ} Helicity from h^{±} Energy Correlations T.E. Coan et al., Physical Review D 55, 7291 (1997)

218. Search for B → μ\overline{ν}_{μ}γ and B → e\overline{ν}_{e}γ T.E. Browder et al., Physical Review D 56, 11 (1997)

219. Study of Gluon versus Quark Fragmentation in Υ → ggγ and e^{+}e^{−} → q\overline{q}γ Events at \sqrt{s} = 10 GeV M.S. Alam et al., Physical Review D 56, 17 (1997)

220. Λ\overline{Λ} Production in Two-photon Interactions S. Anderson et al., Physical Review D 56, R2485 (1997)

221. The Inclusive Decays B → DX and B → D^{∗}X L. Gibbons et al., Physical Review D 56, 3783 (1997)

222. Search for the Decay τ^{−} → 4π^{−}3π^{+}(π^{0})ν_{τ} K.W. Edwards et al., Physical Review D 56, R5297 (1997)

223. Determination of the Michel Parameters and the τ Neutrino Helicity in τ Decay J.P. Alexander et al., Physical Review D 56, 5320 (1997)

224. A New Upper Limit on the Decay η → e^{+}e^{−} T.E. Browder et al., Physical Review D 56, 5359 (1997)

225. Flavor-Specific Inclusive B Decays to Charm T.E. Coan et al., Physical Review Letters 80, 1150 (1998)

226. Search for Inclusive b → sℓ^{+}ℓ^{−} S. Glenn et al., Physical Review Letters 80, 2289 (1998)

227. New Measurement of B → D^{∗}π Branching Fractions G. Brandenburg et al., Physical Review Letters 80, 2762 (1998)

228. Measurement of B(D^{0} → K^{−}π^{+}) using Partial Reconstruction of \overline{B} → D^{∗+}Xℓ^{−}\overline{ν} M. Artuso et al., Physical Review Letters 80, 3193 (1998)

229. Observation of Exclusive Two-Body B Decays to Kaons and Pions R. Godang et al., Physical Review Letters 80, 3456 (1998)

230. Two-Body B Meson Decays to η and η′: Observation of B → η′K B.H. Behrens et al., Physical Review Letters 80, 3710 (1998)

231. Observation of the Radiative Decay D^{∗+} → D^{+}γ J. Bartelt et al., Physical Review Letters 80, 3919 (1998)

232. Investigation of Semileptonic B Meson Decays to p-Wave Charm Mesons A. Anastassov et al., Physical Review Letters 80, 4127 (1998)

233. First Observation of the Cabibbo Suppressed Decay B^{+} → \overline{D}^{0}K^{+} M. Athanas et al., Physical Review Letters 80, 5493 (1998)

234. Observation of B^{+} → ωK^{+} and Search for Related B Decay Modes T. Bergfeld et al., Physical Review Letters 81, 272 (1998)

235. Observation of High Momentum η′ Production in B Decay T.E. Browder et al., Physical Review Letters 81, 1786 (1998)

236. Further Search for the Two-Photon Production of the Glueball Candidate f_{J}(2220) M.S. Alam et al., Physical Review Letters 81, 3328 (1998)

237. First Search for CP Violation in Tau Lepton Decay S. Anderson et al., Physical Review Letters 81, 3823 (1998)

238. A Limit on the Mass of the ν_{τ} R. Ammar et al., Physics Letters B 431, 209 (1998)

239. Measurements of the Meson-Photon Transition Form Factors of Light Pseudoscalar Mesons at Large Momentum Transfer J. Gronberg et al., Physical Review D 57, 33 (1998)

240. Measurement of the Total Cross Section for e^{+}e^{−} → hadrons at \sqrt{s} = 10.52 GeV R. Ammar et al., Physical Review D 57, 1350 (1998)

241. Search for the Decay B → D_{s1}^{+}(2536)X M. Bishai et al., Physical Review D 57, 3847 (1998)

242. Measurement of the Branching Fractions of Λ_{c}^{+} → p\overline{K}n(π)

M.S. Alam et al., Physical Review D 57, 4467 (1998)

243. Search for Color-Suppressed B Hadronic Decay Processes at the Υ(4S) Resonance B. Nemati et al., Physical Review D 57, 5363 (1998)

244. New Limits for Neutrinoless Tau Decays D.W. Bliss et al., Physical Review D 57, 5903 (1998)

245. Study of Semileptonic Decays of B Mesons to Charmed Baryons G. Bonvicini et al., Physical Review D 57, 6604 (1998)

246. Improved Measurement of the Pseudoscalar Decay Constant f_{D_{s}} M. Chadha et al., Physical Review D 58, 032002 (1998)

247. Measurement of the Branching Ratios for the Decays of D_{s}^{+} → ηπ^{+}, η′π^{+}, ηρ^{+}, and η′ρ^{+} C.P. Jessop et al., Physical Review D 58, 052002 (1998)

248. Continuum Charged D^{∗} Spin Alignment at \sqrt{s} = 10.5 GeV G. Brandenburg et al., Physical Review D 58, 052003 (1998)

249. The Hadronic Transition Υ(2S) → Υ(1S)

J.P. Alexander et al., Physical Review D 58, 052004 (1998)

250. Radiative Decay Modes of the D^{0} Meson D. Asner et al., Physical Review D 58, 092001 (1998)

251. First Observation of the Decay τ^{−} → K^{∗−}ην_{τ} M. Bishai et al., Physical Review Letters 82, 281 (1999)

252. First Observation of Υ(1S) → γπ^{+}π^{−} and Υ(1S) → γπ^{0}π^{0} A. Anastassov et al., Physical Review Letters 82, 286 (1999)

253. Observation of Two Narrow States Decaying into Ξ_{c}^{+}γ and Ξ_{c}^{0}γ C.P. Jessop et al., Physical Review Letters 82, 492 (1999)

254. First Observation of the Decay B^{0} → D^{∗+}D^{∗−} M. Artuso et al., Physical Review Letters 82, 3020 (1999)

255. Measurement of the B → Dℓν Branching Fraction and Form Factor J. Bartelt et al., Physical Review Letters 82, 3746 (1999)

256. Measurement of Charm Meson Lifetimes G. Bonvicini et al., Physical Review Letters 82, 4586 (1999)

257. Evidence of New States Decaying into Ξ_{c}^{∗}π J. Alexander et al., Physical Review Letters 83, 3390 (1999)

258. Measurement of the Mass Splittings between the b\overline{b} χ_{b,J}(1P) States K.W. Edwards et al., Physical Review D 59, 032003 (1999)

259. Υ Dipion Transitions at Energies Near the Υ(4S)

S. Glenn et al., Physical Review D 59, 052003 (1999)

260. Search for Baryon and Lepton Number Violating Decays of the τ Lepton R. Godang et al., Physical Review D 59, 091303 (1999)

261. Search for Exclusive Rare Baryonic Decays of B Mesons T.E. Coan et al., Physical Review D 59, 111101 (1999)

262. Study of 3-prong Hadronic τ Decays with Charged Kaons S. Richichi et al., Physical Review D 60, 112002 (1999)

263. Search for Rare and Forbidden η′ Decays R.A. Briere et al., Physical Review Letters 84, 26 (2000)

264. Observation of Radiative Leptonic Decay of the Tau Lepton T. Bergfeld et al., Physical Review Letters 84, 830 (2000)

265. First Observation of the Decay B → J/ψ ϕ K A. Anastassov et al., Physical Review Letters 84, 1393 (2000)

266. Search for the Decay B^{0} → D^{∗0}γ M. Artuso et al., Physical Review Letters 84, 4292 (2000)

267. Search for D^{0}\overline{D}^{0} Mixing R. Godang et al., Physical Review Letters 84, 5038 (2000)

268. Study of Exclusive Radiative B Meson Decays T.E. Coan et al., Physical Review Letters 84, 5283 (2000)

269. Search for CP Violation in B^{±} → J/ψK^{±} and B^{±} → ψ(2S)K^{±} Decays G. Bonvicini et al., Physical Review Letters 84, 5940 (2000)

270. Observation of B → K^{±}π^{0} and B → K^{0}π^{0} and Evidence for B → π^{+}π^{−} D. Cronin-Hennessy et al., Physical Review Letters 85, 515 (2000)

271. Two-body B Meson Decays to η and η′: Observation of B → ηK^{∗} S.J. Richichi et al., Physical Review Letters 85, 520 (2000)

272. Measurement of Charge Asymmetries in Charmless Hadronic B Meson Decays S. Chen et al., Physical Review Letters 85, 525 (2000)

273. Study of Charmless Hadronic B Meson Decays to Pseudoscalar-Vector Final States C. Jessop et al., Physical Review Letters 85, 2881 (2000)

274. Measurements of the Mass, Total Width, and Two-Photon Partial Width of the η_{c} Meson G. Brandenburg et al., Physical Review Letters 85, 3095 (2000)

275. Precise Measurement of B^{0} − \overline{B}^{0} Mixing Parameters at the Υ(4S)

B.H. Behrens et al., Physics Letters B 490, 36 (2000)

276. Hadronic Structure in the Decay τ^{−} → ν_{τ}π^{−}π^{0}π^{0} and the Sign of the Tau Neutrino Helicity D.M. Asner et al., Physical Review D 61, 012002 (2000)

277. Measurement of B → ρℓν Decay and V_{ub} B.H. Behrens et al., Physical Review D 61, 052001 (2000)

278. Limit on the Tau Neutrino Mass from τ^{−} → π^{−}π^{+}π^{−}π^{0}ν_{τ} M. Athanas et al., Physical Review D 61, 052002 (2000)

279. Structure Functions in the Decay τ^{∓} → π^{∓}π^{0}π^{0}ν_{τ} T.E. Browder et al., Physical Review D 61, 052004 (2000)

280. Update of the Search for the Neutrinoless Decay τ → μγ S. Ahmed et al., Physical Review D 61, 071101(R) (2000)

281. Charged Track Multiplicity in B Meson Decay G. Brandenburg et al., Physical Review D 61, 072002 (2000)

282. Resonant Structure of τ → 3ππ^{0}ν_{τ} and τ → ωπν_{τ} Decays K.W. Edwards et al., Physical Review D 61, 072003 (2000)

283. Measurement of the B^{0} and B^{+} Meson Masses from B^{0} → ψ^{(′)}K_{S}^{0} and B^{+} → ψ^{(′)}K^{+} Decays S.E. Csorna et al., Physical Review D 61, 111101(R) (2000)

284. Hadronic Structure in the Decay τ^{−} → π^{−}π^{0}ν_{τ} S. Anderson et al., Physical Review D 61, 112002 (2000)

285. Study of the decays B^{0} → D^{(∗)+}D^{(∗)−} E. Lipeles et al., Physical Review D 62, 032005 (2000)

286. Study of Exclusive Two-Body B^{0} Meson Decays to Charmonium P. Avery et al., Physical Review D 62, 051101(R) (2000)

287. Measurements of Charm Fragmentation into D_{s}^{∗+} and D_{s}^{+} in e^{+}e^{−} Annihilations at \sqrt{s} = 10.5 GeV R.A. Briere et al., Physical Review D 62, 072003 (2000)

288. Measurement of B(Λ_{c}^{+} → pK^{−}π^{+})

D.E. Jaffe et al., Physical Review D 62, 072005 (2000)

289. Resonance Structure of τ^{−} → K^{−}π^{+}π^{−}ν_{τ} Decays D. Asner et al., Physical Review D 62, 072006 (2000)

290. Search for Decays of B^{0} Mesons into Pairs of Leptons: B^{0} → e^{+}e^{−}, B^{0} → μ^{+}μ^{−}, and B^{0} → e^{±}μ^{∓} T. Bergfeld et al., Physical Review D 62, 091102(R) (2000)

291. Measurement of the Product Branching Fraction B(c → Θ_{c}X) × B(Θ_{c} → ΛX) R. Ammar et al., Physical Review D 62, 092007 (2000)

292. Measurement of B → D_{s}^{(∗)+}D^{∗(∗)} Branching Fractions S. Ahmed et al., Physical Review D 62, 112003 (2000)

293. Study of B Decays to Charmonium States B → η_{c}K and B → χ_{c0}K K.W. Edwards et al., Physical Review Letters 86, 30 (2001)

294. First Observation of the Σ_{c}^{∗+} Baryon and a New Measurement of the Σ_{c}^{+} Mass R. Ammar et al., Physical Review Letters 86, 1167 (2001)

295. Measurement of the Λ_{c}^{+} Lifetime A.H. Mahmood et al., Physical Review Letters 86, 2232 (2001)

296. First Observation of the Decays B^{0} → D^{∗−}p\overline{p}π^{+} and B^{0} → D^{∗−}p\overline{n} S. Anderson et al., Physical Review Letters 86, 2732 (2001)

297. Measurement of the Relative Branching Fraction of Υ(4S) to Charged and Neutral B-Meson Pairs J.P. Alexander et al., Physical Review Letters 86, 2737 (2001)

298. Search for B → τν and B → Kν\overline{ν} T.E. Browder et al., Physical Review Letters 86, 2950 (2001)

299. Observation of B → ϕK and B → ϕK^{∗} R.A. Briere et al., Physical Review Letters 86, 3718 (2001)

300. Observation of the Ω_{c}^{0} Charmed Baryon at CLEO D. Cronin-Hennessy et al., Physical Review Letters 86, 3730 (2001)

301. Evidence of New States Decaying into Ξ′_{c}π S.E. Csorna et al., Physical Review Letters 86, 4243 (2001)

302. Study of τ Decay to Six Pions and Neutrino A. Anastassov et al., Physical Review Letters 86, 4467 (2001)

303. Observation of New States Decaying into Λ_{c}^{+}π^{−}π^{+} M. Artuso et al., Physical Review Letters 86, 4479 (2001)

304. Bounds on the CP Asymmetry in Like-Sign Dileptons from B^{0}\overline{B}^{0} Meson Decays D.E. Jaffe et al., Physical Review Letters 86, 5000 (2001)

305. Bounds on the CP Asymmetry in b → sγ Decays T.E. Coan et al., Physical Review Letters 86, 5661 (2001)

306. Experimental Investigation of the Two-Photon Widths of the χ_{c0} and the χ_{c2} Mesons B.I. Eisenstein et al., Physical Review Letters 87, 061801 (2001)

307. Rate Measurement of D^{0} → K^{−}π^{+}π^{0} and Constraints on D^{0} - \overline{D}^{0} Mixing G. Brandenburg et al., Physical Review Letters 87, 071802 (2001)

308. Search for the Decay Υ(1S) → γη′ S.J. Richichi et al., Physical Review Letters 87, 141801 (2001)

309. Improved Upper Limits on the Flavor-Changing Neutral Current Decays B → Kℓ^{+}ℓ^{−} and B → K^{∗}(892) ℓ^{+}ℓ^{−} S. Anderson et al., Physical Review Letters 87, 181803 (2001)

310. First Measurement of Γ(D^{∗+})

S. Ahmed et al., Physical Review Letters 87, 251801 (2001)

311. Branching Fraction and Photon Energy Spectrum for b → sγ S. Chen et al., Physical Review Letters 87, 251807 (2001)

312. Hadronic Mass Moments in Inclusive Semileptonic B Meson Decays D. Cronin-Hennessy et al., Physical Review Letters 87, 251808 (2001)

313. Search for the Familon via B^{±} → π^{±}X^{0}, B^{±} → K^{±}X^{0}, and B^{0} → K_{S}^{0}X^{0} Decays R. Ammar et al., Physical Review Letters 87, 271801 (2001)

314. Study of χ_{c1} and χ_{c2} Production in B Meson Decays S. Chen et al., Physical Review D 63, 031102(R) (2001)

315. Study of B → ψ(2S)K and B → ψ(2S)K^{∗}(892) Decays S. Richichi et al., Physical Review D 63, 031103(R) (2001)

316. Search for a Scalar Bottom Quark with Mass 3.5–4.5 GeV/c^{2} V. Savinov et al., Physical Review D 63, 051101(R) (2001)

317. Search for CP Violation in D^{0} → K_{S}^{0}π^{0}, D^{0} → π^{0}π^{0} and D^{0} → K_{S}^{0}K_{S}^{0} G. Bonvicini et al., Physical Review D 63, 071101(R) (2001)

318. Dalitz Analysis of the Decay D^{0} → K^{−}π^{+}π^{0} S. Kopp et al., Physical Review D 63, 092001 (2001)

319. A Measurement of the Decay Asymmetry Parameters in Ξ_{c}^{0} → Ξ^{−}π^{+} S. Chen et al., Physical Review D 63, 111102(R) (2001)

320. Correlated Λ_{c}^{+}\overline{Λ}_{c}^{−} production in e^{+}e^{−} Annihilations at \sqrt{s} ∼ 10.5 GeV A. Bornheim et al., Physical Review D 63, 112003 (2001)

321. Search for the Decay B^{+} → D^{∗+}K_{S}^{0} A. Gritsan et al., Physical Review D 64, 077501 (2001)

322. First Observation of B → D^{(∗)}ρ^{′−}, ρ^{′−} → ωπ^{−} J.P. Alexander et al., Physical Review D 64, 092001 (2001)

323. Search for CP Violation in τ → ππ^{0}ν_{τ} Decay P. Avery et al., Physical Review D 64, 092005 (2001)

324. Evidence for the Decay D^{0} → K^{+}π^{−}π^{+}π^{−} S.A. Dytman et al., Physical Review D 64, 111101(R) (2001)

325. A Search for Charmless B → V V Decay R. Godang et al., Physical Review Letters 88, 021802 (2002), hep-ex/0101029

326. Observation of \overline{B}^{0} → D^{0}π^{0} and \overline{B}^{0} → D^{∗0}π^{0} T.E. Coan et al., Physical Review Letters 88, 062001 (2002), hep-ex/0110055

327. Observation of Exclusive \overline{B} → D^{(∗)}K^{∗−} Decays R. Mahapatra et al., Physical Review Letters 88, 101803 (2002), hep-ex/0112033

328. Search for CP Violation in τ → Kπν_{τ} Decays G. Bonvicini et al., Physical Review Letters 88, 111803 (2002), hep-ex/0111095

329. Improved Measurement of |V_{ub}| with Inclusive Semileptonic B Decays A. Bornheim et al., Physical Review Letters 88, 231803 (2002), hep-ex/0202019

330. Improved Measurement of |V_{cb}| using \overline{B} → D^{∗}ℓ\overline{ν} Decays R.A. Briere et al., Physical Review Letters 89, 081803 (2002), hep-ex/0203032

331. Observation of the Decay Ω_{c}^{0} → Ω^{−}e^{+}ν_{e} R. Ammar et al., Physical Review Letters 89, 171803 (2002), hep-ex/0207078

332. Measurement of D^{+} → K^{∗0}ℓ^{+}ν_{ℓ} Branching Fraction G. Brandenburg et al., Physical Review Letters 89, 222001 (2002), hep-ex/0203030

333. Observation of B → K_{S}^{0}π^{+}π^{−} and Evidence for B → K^{∗±}π^{∓} E. Eckhart et al., Physical Review Letters 89, 251801 (2002), hep-ex/0206024

334. Dalitz Analysis of D^{0} → K_{S}^{0}π^{+}π^{−} H. Muramatsu et al., Physical Review Letters 89, 251802 (2002), hep-ex/0207067

335. Measurements of Inclusive B → ψ Production S. Anderson et al., Physical Review Letters 89, 282001 (2002), hep-ex/0207059

336. First Observation of \overline{B}^{0} → D^{∗0}π^{+}π^{+}π^{−}π^{−} K.W. Edwards et al., Physical Review D 65, 012002 (2002), hep-ex/0105071 337. Measurement of the Ξ_{c}^{+} Lifetime

A.H. Mahmood et al., Physical Review D 65, 031102(R) (2002), hep-ex/0110058

338. Search for B^{0} → π^{0}π^{0} Decay D.M. Asner et al., Physical Review D 65, 031103(R) (2002), hep-ex/0103040

339. First Measurement of Γ(D^{∗+}) and Precision Measurement of m_{D^{∗+}} − m_{D^{0}} A. Anastassov et al., Physical Review D 65, 032003 (2002), hep-ex/0108043

340. Measurement of the Masses and Widths of the Σ_{c}^{++} and Σ_{c}^{0} Charmed Baryons M. Artuso et al., Physical Review D 65, 071101(R) (2002), hep-ex/0110071

341. Further Experimental Studies of Two-Body Radiative Upsilon Decays G. Masek et al., Physical Review D 65, 072002 (2002), hep-ex/0201005

342. Lifetime Differences, Direct CP Violation, and Partial Widths in D^{0} Meson Decays to K^{+}K^{−} and π^{+}π^{−} S.E. Csorna et al., Physical Review D 65, 092001 (2002), hep-ex/0111024

343. Search for Lepton-Flavor-Violating Decays of B Mesons K.W. Edwards et al., Physical Review D 65, 111102(R) (2002), hep-ex/0408011

344. Measurement of B(B^{−} → D^{0}π^{−}) and B(B^{0} → D^{+}π^{−}) and Isospin Analysis of B → Dπ Decays S. Ahmed et al., Physical Review D 66, 031101(R) (2002), hep-ex/0206030

345. Correlated Inclusive Λ\overline{Λ} Production in e^{+}e^{−} Annihilations at \sqrt{s} ∼ 10.5 GeV Z. Metreveli et al., Physical Review D 66, 052002 (2002)

346. Measurement of the Ratio of Branching Fractions of the Υ(4S) to Charged and Neutral B Mesons S.B. Athar et al., Physical Review D 66, 052003 (2002), hep-ex/0202033

347. Search for Neutrinoless tau Decays Involving the K_{S}^{0} Meson S. Chen et al., Physical Review D 66, 071101 (2002), hep-ex/0208019

348. Antisearch for the Glueball Candidate f_{J}(2220) in Two-Photon Interactions K. Benslama et al., Physical Review D 66, 077101 (2002), hep-ex/0204019

349. Measurement of Exclusive B Decays to Final States Containing a Charmed Baryon S.A. Dytman et al., Physical Review D 66, 091101(R) (2002), hep-ex/0208006

350. First Search for the Flavor Changing Neutral Current Decay D^{0} → γγ T.E. Coan et al., Physical Review Letters 90, 101801 (2003), hep-ex/0212045

351. Branching Fractions of tau Leptons to Three Charged Hadrons R.A. Briere et al., Physical Review Letters 90, 181802 (2003), hep-ex/0302028

352. First Observation of the Exclusive Decays Λ_{c}^{+} → Λπ^{+}π^{+}π^{−}π^{0} and Λ_{c}^{+} → Λωπ^{+} D. Cronin-Hennessy et al., Physical Review D 67, 012001 (2003), hep-ex/0210048 353. Determination of the \overline{B} → D^{∗}ℓ\overline{ν} Decay Width and |V_{cb}|

N.E. Adam et al., Physical Review D 67, 032001 (2003)

354. Inclusive η′ Production at the Υ(1S)

M. Artuso et al., Physical Review D 67, 052003 (2003), hep-ex/0211029

355. Measurement of Lepton Momentum Moments in the Decay \overline{B} → Xℓ\overline{ν} and Determination of Heavy Quark Expansion Parameters and |V_{cb}| A. H. Mahmood et al., Physical Review D 67, 072001 (2003), hep-ex/0212051

356. Measurements of the Branching Fractions and Helicity Amplitudes in B → D^{∗}ρ Decays S.E. Csorna et al., Physical Review D 67, 112002 (2003), hep-ex/0301028

357. Measurement of the Charge Asymmetry in B → K^{∗}(892)^{±}π^{∓} B. Eisenstein et al., Physical Review D 68, 017101 (2003), hep-ex/0304036

358. Study of the Charmless Inclusive B → η′X Decay G. Bonvicini et al., Physical Review D 68, 011101(R) (2003), hep-ex/0303009

359. Search for B → pe^{−}\overline{ν}_{e}X Decay Using a Partial Reconstruction Method N.E. Adam et al., Physical Review D 68, 012004 (2003), hep-ex/0304015

360. Search for Baryons in the Radiative Penguin Decay b → sγ K.W. Edwards et al., Physical Review D 68, 011102(R) (2003), hep-ex/0305005

361. Observation of a Narrow resonance of Mass 2.46 GeV/c^{2} Decaying to D_{s}^{∗+}π^{0} and Confirmation of the D_{sJ}^{∗}(2317) State D. Besson et al., Physical Review D 68, 032002 (2003), hep-ex/0305100

362. Measurements of Charmless Hadronic Two-Body B Meson Decays and the Ratio B(B → DK)/B(B → Dπ) A. Bornheim et al., Physical Review D 68, 052002 (2003), hep-ex/0302026

363. Study of the q^{2}-dependence of B → πℓν and B → ρ(ω)ℓν Decay and Extraction of |V_{ub}| S.B. Athar et al., Physical Review D 68, 072003 (2003), hep-ex/0304019

364. Observation of η′_{c} Production in γγ Fusion at CLEO D. Asner et al., Physical Review Letters 92, 142001 (2004), hep-ex/0312058

365. Observation of the Hadronic Transitions χ_{b1,2}(2P) → ωΥ(1S)

D. Cronin-Hennessy et al., Physical Review Letters 92, 222002 (2004), hep-ex/0311043

366. Wess-Zumino Current and the Structure of the Decay τ → K^{−}K^{+}π^{−}ν_{τ} T.E. Coan et al., Physical Review Letters 92, 232001 (2004), hep-ex/0401005

367. First Observation and Dalitz Analysis of the D^{0} → K_{S}^{0}ηπ^{0} Decay P. Rubin et al., Physical Review Letters 93, 111801 (2004), hep-ex/0405011 368. Search for the Lepton-Flavor-Violating Leptonic B Decays B^{0} → μ^{±}τ^{∓} and

B^{0} → e^{±}τ^{∓} A. Bornheim et al., Physical Review Letters 93, 241802 (2004), hep-ex/0408011

369. Measurement of the Decay Rate of Ξ_{c}^{0} → pK^{−}K^{−}π^{+} Relative to Ξ_{c}^{0} → Ξ^{−}π^{+} I. Danko et al., Physical Review D 69, 052004 (2004), hep-ex/0309020

370. Cabibbo-Suppressed Decays of D^{+} → π^{+}π^{0}, K^{+}\overline{K}^{0}, K^{+}π^{0} K. Arms et al., Physical Review D 69, 071102(R) (2004), hep-ex/0309065

371. First Observation of a Υ(1D) State G. Bonvicini et al., Physical Review D 70, 032001 (2004), hep-ex/0404021

372. Moments of the B Meson Inclusive Semileptonic Decay Rate using Neutrino Reconstruction S.E. Csorna et al., Physical Review D 70, 032002 (2004), hep-ex/0403052

373. Measurement of the B-Meson Inclusive Semileptonic Branching Fraction and Electron-Energy Moments A.H. Mahmood et al., Physical Review D 70, 032003 (2004), hep-ex/0403053

374. New Measurements of Υ(1S) Decays to Charmonium Final States R.A. Briere et al., Physical Review D 70, 072001 (2004), hep-ex/0407030

375. Search for CP Violation in D^{0} → K_{S}^{0}π^{+}π^{−} D. Asner et al., Physical Review D 70, 091101(R) (2004), hep-ex/0311033

376. Charm Meson Spectra in e^{+}e^{−} Annihilation at 10.5 GeV C.M.E.

M. Artuso et al., Physical Review D 70, 112001 (2004), hep-ex/0402040

377. Photon Transitions in ψ(2S) Decays to χ_{cJ}(1P) and η_{c}(1S)

S.B. Athar et al., Physical Review D 70, 112002 (2004), hep-ex/0408133

378. Measuring B(D^{+} → μ^{+}ν) and the Pseudoscalar Decay Constant f_{D^{+}} G. Bonvicini et al., Physical Review D 70, 112004 (2004), hep-ex/0411050

379. Study of the Semileptonic Charm Decays D^{0} → π^{−}ℓ^{+}ν and D^{0} → K^{−}ℓ^{+}ν G.S. Huang et al., Physical Review Letters 94, 011802 (2005), hep-ex/0407035

380. Measurement of the Muonic Branching Fractions of the Narrow Upsilon Resonances G.S. Adams et al., Physical Review Letters 94, 012001 (2005), hep-ex/0409027

381. Observation of 1^{−}0^{−} Final States from ψ(2S) Decays and e^{+}e^{−} Annihilation N.E. Adam et al., Physical Review Letters 94, 012005 (2005), hep-ex/0407028

382. Photon Transitions in Υ(2S) and Υ(3S) Decays M. Artuso et al., Physical Review Letters 94, 032001 (2005), hep-ex/0411068

383. Search for X(3872) in γγ Fusion and ISR at CLEO S. Dobbs et al., Physical Review Letters 94, 032004 (2005), hep-ex/0410038 384. Improved Measurement of the Form Factors in the Decay Λ_{c}^{+} → Λe^{+}ν_{e}

J. W. Hinson et al., Physical Review Letters 94, 191801 (2005), hep-ex/0501002

385. Branching Fractions for ψ(2S) to J/ψ Transitions N.E. Adam et al., Physical Review Letters 94, 232002 (2005), hep-ex/0503028

386. Study of τ Decays to Four-Hadron Final States with Kaons K. Arms et al., Physical Review Letters 94, 241802 (2005), hep-ex/0501042

387. Observation of Thirteen New Exclusive Multi-Body Hadronic Decays of the ψ(2S) R.A. Briere et al., Physical Review Letters 95, 062001 (2005), hep-ex/0505101

388. Observation of the h_{c}(^{1}P_{1}) State of Charmonium J. L. Rosner et al., Physical Review Letters 95, 102003 (2005), hep-ex/0505073

389. Measurement of Absolute Hadronic Branching Fractions of D Mesons and e^{+}e^{−} → D\overline{D} Cross Sections at E_{cm} = 3773 MeV Q. He et al., Physical Review Letters 95, 121801 (2005), hep-ex/0504003 Erratum Physical Review Letters 96, 199903 (2006)

390. Absolute Branching Fraction Measurements of Exclusive D^{+} Semileptonic Decays G. S. Huang et al., Physical Review Letters 95, 181801 (2005), hep-ex/0506053

391. Absolute Branching Fraction Measurements of Exclusive D^{0} Semileptonic Decays T.E. Coan et al., Physical Review Letters 95, 181802 (2005), hep-ex/0506052

392. Search for Rare and Forbidden Decays D^{+} → h^{±}e^{∓}e^{+} Q. He et al., Physical Review Letters 95, 221802 (2005), hep-ex/0508031

393. Improved Measurement of B(D^{+} → μ^{+}ν) and the Pseudoscalar Decay Constant f_{D^{+}} M. Artuso et al., Physical Review Letters 95, 251801 (2005), hep-ex/0508057

394. Evidence for B_{s}^{(∗)}\overline{B}_{s}^{(∗)} Production at the Υ(5S)

M. Artuso et al., Physical Review Letters 95, 261801 (2005), hep-ex/0508047

395. Precision Measurements of the Timelike Electromagnetic Form Factors of Pion, Kaon, and Proton T.K. Pedlar et al., Physical Review Letters 95, 261803 (2005), hep-ex/0510005

396. Search for e^{+}e^{−} → Λ_{b}^{0}\overline{Λ}_{b}^{0} Near Threshold D. Besson et al., Physical Review D 71, 012004 (2005), hep-ex/0411078

397. A New Measurement of the Masses and Widths of the Σ_{c}^{∗++} and Σ_{c}^{∗0} Charmed Baryons S.B. Athar et al., Physical Review D 71, 051101(R) (2005), hep-ex/0410088

398. Limits on Neutral D Mixing in Semileptonic Decays C. Cawlfield et al., Physical Review D 71, 077101 (2005), hep-ex/0502012 399. The Search for η(1440) → K_{S}^{0}K^{±}π^{∓} in Two-Photon Fusion at CLEO

R. Ahohe et al., Physical Review D 71, 072001 (2005), hep-ex/0501026

400. Measurement of the Branching Fractions for J/ψ → ℓ^{+}ℓ^{−} Z. Li et al., Physical Review D 71, 111103(R) (2005), hep-ex/0503027

401. Search for D^{0}-\overline{D}^{0} Mixing in the Dalitz Plot Analysis of D^{0} → K_{S}^{0}π^{+}π^{−} D.M. Asner et al., Physical Review D 72, 012001 (2005), hep-ex/0503045

402. Searches for CP Violation and ππ S-Wave in the Dalitz-Plot Analysis of D^{0} → π^{+}π^{−}π^{0} D. Cronin-Hennessy et al., Physical Review D 72, 031102(R) (2005), hep-ex/0503052

403. Branching Fraction Measurements of ψ(2S) Decay to Baryon-Antibaryon Final States T.K. Pedlar et al., Physical Review D 72, 051108(R) (2005), hep-ex/0505057

404. Observation of the ^{1}P_{1} State of Charmonium P. Rubin et al., Physical Review D 72, 092004 (2005), hep-ex/0508037

405. Observation of B_{s} Production at the Υ(5S) Resonance G. Bonvicini et al., Physical Review Letters 96, 022002 (2006), hep-ex/0510034

406. Search for Exclusive Multi-Body Non-D\overline{D} Decays at the ψ(3770)

G.S. Huang et al., Physical Review Letters 96, 032003 (2006), hep-ex/0509046

407. New Measurements of Cabibbo-Suppressed Decays of D Mesons with the CLEO-c Detector P. Rubin et al., Physical Review Letters 96, 081802 (2006), hep-ex/0512063

408. Observation of ψ(3770) → ππJ/ψ and Measurement of Γ_{ee}[ψ(2S)] N. E. Adam et al., Physical Review Letters 96, 082004 (2006), hep-ex/0508023

409. Measurement of σ(e^{+}e^{−} → ψ(3770) → hadrons) at E_{cm}=3773 MeV D. Besson et al., Physical Review Letters 96, 092002 (2006), hep-ex/0512038

410. Dielectron Widths of the Υ(1S, 2S, 3S) Resonances J. L. Rosner et al., Physical Review Letters 96, 092003 (2006), hep-ex/0512056

411. Experimental Limits on Weak Annihilation Contributions to b → uℓν Decay J.L. Rosner et al., Physical Review Letters 96, 121801 (2006), hep-ex/0601027

412. Measurements of the Exclusive Decays of the Υ(5S) to B Meson Final States and Improved B_{s}^{∗} Mass Measurement O. Aquines et al., Physical Review Letters 96, 152001 (2006), hep-ex/0601044

413. Charmonium Decays of Y (4260), ψ(4160), and ψ(4040)

T.E. Coan et al., Physical Review Letters 96, 162003 (2006), hep-ex/0602034

414. Observation of ψ(3770) → γχ_{c1} → γγJ/ψ T.E. Coan et al., Physical Review Letters 96, 182002 (2006), hep-ex/0509030 415. Absolute Branching Fraction Measurements for D^{+} and D^{0} Inclusive Semileptonic

Decays N.E. Adam et al., Physical Review Letters 97, 251801 (2006), hep-ex/0604044

416. Decay of ψ(3770) to Light Hadrons G.S. Adams et al., Physical Review D 73, 012002 (2006), hep-ex/0509011

417. Experimental Study of χ_{b}(2P) → ππχ_{b}(1P)

C. Cawlfield et al., Physical Review D 73, 012003 (2006), hep-ex/0511019

418. Radiative Decays of the Υ(1S) to a Pair of Charged Hadrons S.B. Athar et al., Physical Review D 73, 032001 (2006), hep-ex/0510015

419. Measurement of Γ_{ee}(J/ψ), Γ_{tot}(J/ψ), and Γ_{ee}[ψ(2S)]/Γ_{ee}(J/ψ)

G.S. Adams et al., Physical Review D 73, 051103(R) (2006), hep-ex/0512046

420. Two-Photon Width of the Charmonium State χ_{c2} S. Dobbs et al., Physical Review D 73, 071101(R) (2006), hep-ex/0510033

421. An Investigation of D^{+} → τ^{+}ν P. Rubin et al., Physical Review D 73, 112005 (2006), hep-ex/0604043

422. Measurement of the Direct Photon Momentum Spectrum in Υ(1S), Υ(2S), and Υ(3S) Decays D. Besson et al., Physical Review D 74, 012003 (2006), hep-ex/0512061

423. Measurement of Interference between Electromagnetic and Strong Amplitudes in ψ(2S) Decays to Two Pseudoscalar Mesons S. Dobbs et al., Physical Review D 74, 011105(R) (2006), hep-ex/0603020

424. Search for the non-D\overline{D} Decay ψ(3770) → K_{S}^{0}K_{L}^{0} D. Cronin-Hennessy et al., Physical Review D 74, 012005 (2006), hep-ex/0603026

425. Observation of ψ(3770) → γχ_{c0} R.A. Briere et al., Physical Review D 74, 031106(R) (2006), hep-ex/0605070

426. Measurement of Interfering K^{∗+}K^{−} and K^{∗−}K^{+} Amplitudes in the Decay D^{0} → K^{+}K^{−}π^{0} C. Cawlfield et al., Physical Review D 74, 031108(R) (2006), hep-ex/0606045

427. Model Independent Measurement of Form Factors in the Decay D^{+} → K^{−}π^{+}e^{+}ν_{e} M.R. Shepherd et al., Physical Review D 74, 052001 (2006), hep-ex/0606010

428. Branching Fraction for the Doubly-Cabibbo-Suppressed Decay D^{+} → K^{+}π^{0} S.A. Dytman et al., Physical Review D 74, 071102(R) (2006), hep-ex/0609008 Erratum Physical Review D 74, 079904 (2006)

429. Confirmation of the Y (4260) Resonance Production in ISR Q. He et al., Physical Review D 74, 091104(R) (2006), hep-ex/0611026 430. Improved Measurement of the Branching Fraction and Energy Spectrum of η′ from

Υ(1S) Decays O. Aquines et al., Physical Review D 74, 092006 (2006), hep-ex/0610032

431. Measurement of Inclusive Production of η, η′ and ϕ Mesons in D^{0}, D^{+} and D_{s}^{+} Decays G.S. Huang et al., Physical Review D 74, 112005 (2006), hep-ex/0610008

432. Observation of Υ(3S) → ττ and Tests of Lepton Universality in Υ Decays D. Besson et al., Physical Review Letters 98, 052002 (2007), hep-ex/0607019

433. Precision Determination of the D^{0} Mass C. Cawlfield et al., Physical Review Letters 98, 092002 (2007), hep-ex/0701016

434. Study of Exclusive Charmless Semileptonic B Decay and |V_{ub}| N.E. Adam et al., Physical Review Letters 99, 041802 (2007), hep-ex/0703041

435. Measurement of the Decay Constant f_{D}+ s using D_{s}^{+} → ℓ^{+}ν M. Artuso et al., Physical Review Letters 99, 071802 (2007), arXiv:0704.0629

436. Measurement of Prominent η Decay Branching Fractions A. Lopez et al., Physical Review Letters 99, 122001 (2007), arXiv:0707.1601

437. Measurement of the η-Meson Mass Using ψ(2S) → ηJ/ψ D. Miller et al., Physical Review Letters 99, 122002 (2007), arXiv:0707.1810

438. Evidence for the Decay D^{0} → K^{−}π^{+}π^{−}e^{+}ν_{e} M. Artuso et al., Physical Review Letters 99, 191801 (2007), arXiv:0705.4276

439. Suppressed Decays of D_{s}^{+} Mesons to Two Pseudoscalar Mesons G.S. Adam et al., Physical Review Letters 99, 191805 (2007), arXiv:0708.0139

440. Measurement of B(Υ(5S) → B_{s}^{(∗)}\overline{B}_{s}^{(∗)}) using ϕ Mesons G.S. Huang et al., Physical Review D 75, 012002 (2007), hep-ex/0610035

441. Search for ψ(2s) → η_{c}π^{+}π^{−}π^{0} T. Pedlar et al., Physical Review D 75, 011102(R) (2007), hep-ex/0611027

442. Antideuteron Produciton in Υ(nS) Decays and the Nearby Continuum D.M. Asner et al., Physical Review D 75, 012009 (2007), hep-ex/0612019

443. χ_{cJ} Decays to h^{+}h^{−}h^{0} S.B. Athar et al., Physical Review D 75, 032002 (2007), hep-ex/0611032

444. Search for Invisible Decays of the Υ(1S) Resonance P. Rubin et al., Physical Review D 75, 031104(R) (2007), hep-ex/0612051

445. χ_{c0} and χ_{c2} Decays into ηη, ηη′, and η′η′ Final States G.S. Adams et al., Physical Review D 75, 071101(R) (2007), hep-ex/0611021 446. Radiative Decays of the Υ(1S) → γπ^{0}π^{0}, γηη and γπ^{0}η

D. Besson et al., Physical Review D 75, 072001 (2007), hep-ex/0512003

447. Dalitz Plot Analysis of the D^{+} → π^{−}π^{+}π^{+} Decay G. Bonvicini et al., Physical Review D 76, 012001 (2007), arXiv:0704.3954

448. Comparison of Particle Production in Quark and Gluon Fragmentation at \sqrt{s} ∼ 10 GeV R.A. Briere et al., Physical Review D 76, 012005 (2007), arXiv:0704.2766

449. Study of Exclusive Charmless Semileptonic B Decays and Extraction of |V_{ub}| at CLEO D. M. Asner et al., Physical Review D 76, 012007 (2007), hep-ex/0703042

450. Study of Dipion Transitions Among Upsilon(3S), Upsilon(2S) and Upsilon(1S) States D. Cronin-Hennessy et al., Physical Review D 76, 072001 (2007), arXiv:0706.2317

451. Measurement of B(D_{s}^{+} → ℓ^{+}ν) and the Decay Constant f_{D}+ s T. K. Pedlar et al., Physical Review D 76, 072002 (2007), arXiv:0704.0437

452. Search for Radiative Decays of Υ(1S) into η and η′ S. B. Athar et al., Physical Review D 76, 072003 (2007), arXiv:0704.3063

453. Measurement of the Total Hadronic Cross Section in e^{+}e^{−} Annihilations Below 10.56 GeV D. Besson et al., Physical Review D 76, 072008 (2007), arXiv:0706.2813

454. Measurement of Upper Limits for Υ → γ + R Decays J.L. Rosner et al., Physical Review D 76, 117102 (2007), arXiv:0704.2773

455. Measurement of Absolute Hadronic Branching Fractions of D Mesons and e^{+}e^{−} → D\overline{D} Cross Sections at the ψ(3770) S. Dobbs et al., Physical Review D 76, 112001 (2007), arXiv:0709.3783

456. Comparison of D → K_{S}^{0}π and D → K_{L}^{0}π Decay Rates Q. He et al., Physical Review Letters 100, 091801 (2008), arXiv:0711.1463

457. A Study of the Decays D^{0} → π^{−}e^{+}ν_{e}, D^{0} → K^{−}e^{+}ν_{e}, D^{+} → π^{0}e^{+}ν_{e} and D^{+} → \overline{K}^{0}e^{+}ν_{e} D. Cronin-Hennessy et al., CLNS 06/1967, CLEO 06–12 (submitted to Physical Review Letters), arXiv:0712.0998

458. A Study of the Semileptonic Charm Decays D^{0} → π^{−}e^{+}ν_{e}, D^{+} → π^{0}e^{+}ν_{e}, D^{0} → K^{−}e^{+}ν_{e}, and D^{+} → \overline{K}^{0}e^{+}ν_{e} S. Dobbs et al., CLNS 06/1968, CLEO 06–13 (submitted to Physical Review D), arXiv:0712.1020

459. Observation of D^{+} → ηe^{+}ν_{e} R. E. Mitchell et al., CLNS 07/2010, CLEO 07–14 (submitted to Physical Review Letters), arXiv:0802.4222 460. Measurement of the Absolute Branching Fraction of D_{s}^{+} → τ^{+}ν_{τ} Decay

K.M. Ecklund et al., CLNS 07/2011, CLEO 07–15 (submitted to Physical Review Letters), arXiv:0712.1175

461. Absolute Branching Fractions of Cabibbo-Suppressed D → K\overline{K} Decays G. Bonvicini et al., CLNS 07/2012, CLEO 07–16 (submitted to Physical Review D), arXiv:0803.0793

462. Determination of the Strong Phase in D^{0} → K^{+}π^{−} Using Quantum-Correlated Measurements J. Rosner et al., CLNS 07/2013, CLEO 07–17 (submitted to Physical Review Letters), arXiv:0802.2264

463. Determination of the D^{0} → K^{+}π^{−} Relative Strong Phase Using Quantum-Correlated Measurements in e^{+}e^{−} → D^{0}\overline{D}^{0} at CLEO D. Asner et al., CLNS 07/2014, CLEO 07–18 (submitted to Physical Review D), arXiv:0802.2268

464. Measurement of Charm Production Cross Sections in e^{+}e^{−} Annihilation at Energies between 3.97 and 4.26 GeV D. Cronin-Hennessy et al., CLNS 07/2015, CLEO 07–19 (submitted to Physical Review D), arXiv:0801.3418

465. Absolute Measurement of Hadronic Branching Fractions of the D_{s}^{+} Meson J.P. Alexander et al., CLNS 07/2016, CLEO 07–20 (submitted to Physical Review Letters), arXiv:0801.0680

466. Measurement of Exclusive D Meson Decays to η and η′ Final States and SU(3) Amplitude Analysis M. Artuso et al., CLNS 08/2017, CLEO 08–01 (submitted to Physical Review D), arXiv:0802.2664

467. Dalitz Plot Analysis of the D^{+} → K^{−}π^{+}π^{+} Decay G. Bonvicini et al., CLNS 08/2018, CLEO 08–02 (submitted to Physical Review D), arXiv:0802.4214

468. First Observation of the Deacy D_{s}^{+} → p\overline{n} S. Athar et al., CLNS 08/2022, CLEO 08–06 (submitted to Physical Review Letters), arXiv:0803.1118

469. Two-Photon Widths of the χ_{cJ} States of Charmonium K.M. Ecklund et al., CLNS 08/2019, CLEO 08–03 (submitted to Physical Review Letters), arXiv:0803.2869